# A note on division rings satisfying generalized rational identities with anti-automorphisms

### v1

@@ 1. Introduction and the main result
A classic result states that if $D$ satisfies the polynomial identity, then $D$ is centrally finite (see [9]). To present this result and the extended results we aim to prove, we recall some concepts. Let $X=\{x_{1},x_{2},\cdots,x_{n}\}$ be a set of $n$ non-commutative indeterminates and $K$ be a field. We denote $K\langle X\rangle$ as the polynomial ring in indeterminates $X$ over $K$, and $R$ as a ring whose center contains $K$. A polynomial $f(X)\in K\langle X\rangle\backslash\{0\}$ is called a polynomial identity of $R$ if for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in R^{n}$, we have $f(a_{1},a_{2},\dots,a_{n})=0$. In this case, we say $R$ satisfies the polynomial identity $f$. If $R$ satisfies some polynomial identity, then we say $R$ satisfies a polynomial identity. Since we only work with division rings, we directly assume $R=D$ is a division ring and $K$ coincides with the center $F$ of $D$.
@@
The most classical result concerning division rings satisfying polynomial identities relates to its finite-dimensionality. Specifically, it is proven in [9] that if $D$ is a division ring with center $F$ and satisfies some polynomial identity over $F$, then $D$ is centrally finite. This result has been subsequently extended in many different directions (e.g., see [1, 2, 3, 5, 12, 10, 11]). More precisely, we denote $D\langle X\rangle=D*_{F}F\langle X\rangle$ as the free product of $D$ with $F\langle X\rangle$ over $F$. Each element in $D\langle X\rangle\backslash\{0\}$ is called a generalized polynomial over $D$. Furthermore, since the ring of generalized polynomials $D\langle X\rangle$ is fir, the maximal right ring of quotients of $D\langle X\rangle$ is a division ring (e.g., see [6]). This division ring is denoted as $D(X)$. Each nonzero element in $D(X)$ is called a generalized rational polynomial over $D$. Suppose $f(X)$ is a generalized rational polynomial (respectively, a generalized polynomial) over $D$. If for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in D^{n}$ such that when $f(a_{1},a_{2},\dots,a_{n})$ is defined, one has $f(a_{1},a_{2},\dots,a_{n})=0$, then we say $D$ satisfies the generalized rational identity (respectively, the generalized identity) $f$, or $f$ is a generalized rational identity (respectively, a generalized identity) of $D$.
@@
In 1966, Amitsur proved that if the center $F$ is infinite and the division ring $D$ satisfies a generalized rational identity, then $D$ is centrally finite (see [2]). There are several extensions of this result. For example, in 1996, Chiba proved that if the center $F$ is infinite and the multiplicative group $D^{*}=D\backslash\{0\}$ contains a non-central proper subnormal subgroup satisfying a generalized rational identity, then $D$ is centrally finite (see [4]). Another direction is to consider identities corresponding to an anti-automorphism. Recall that a bijection $\sigma:D\to D,x\mapsto x^{\sigma},$ is called an anti-automorphism of $D$ if $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma}$ and $(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ for all $a,b\in D$. In the case where the order of $\sigma$ is $2$, $\sigma$ is called an involution of $D$. Let $m$ be a positive integer such that $\sigma^{i}\neq\mathrm{Id}$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, denote $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ as the set of indeterminates indexed corresponding to each $\sigma^{i}$ and $X_{m}=\cup_{i=0}^{m}X^{\sigma^{i}}$. We denote $D(X_{m})$ as the division ring as mentioned above. Let $f(X_{m})$ be a nonzero element in $D(X_{m})$. If for every $n$-tuple $(r_{1},r_{2},\cdots,r_{n})\in D^{n}$ such that when
@@
then we say $f(X_{m})$ is a generalized rational identity of $D$ with respect to $\sigma^{m}$. In this case, we simply say that $D$ satisfies a $\sigma^{m}$- GRI $f$ or $f$ is a $\sigma^{m}$-GRI of $D$. The ”best” result is related to when $\sigma$ is an involution. In this case, $m=1$ and $\sigma$ will be denoted as $\star$. In more detail, it is proved that if the center $F$ is infinite and $D$ satisfies a $\star$-GRI, then $D$ is centrally finite (see [7]). The case of generalized polynomial identities with respect $\sigma$ receives even more attention (e.g., see the book [3] for a survey on this topic). The objective of this paper is to prove the following result.
@@ 2. Preliminaries
In short and simple terms, we prove the main result of this paper by induction on $n$, the number of indeterminates in the identity. In this section, we prove the case for $n=1$. To do it, we need to review some basic concepts and establish notation. Many results in this section may have proofs similar to those found in some reference documents, but none of the results are directly quoted from them. Therefore, for the convenience of the readers, we will provide proofs for the results if necessary. The main result we obtained in this section is Lemma  2.4, which states that if a division ring satisfies a homogeneous linear identity in one variable, then this division ring is centrally finite.
@@
We also define the $\sigma^{m}$-height of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}\text{-}ht_{x_{j}}(P)$, as the maximum value between $(\sigma^{m}\text{-}\deg_{x_{j}}(P)-1)$ and 00. Similarly, the height of $P$, denoted $ht(P)$, is defined as
@@
In other words, the height of $P$ is equal to the degree of $P$ minus the number of distinct indeterminates appearing in $P$. Let $\AA$ be a $F$-basis of $D$ containing $1$. Then $\AA$, combined with the standard $F$-basis of $F\langle X_{m}\rangle$, induces a monomial basis $M(\AA)$ for $D\langle X_{m}\rangle$. Suppose $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ where $P_{1},P_{2},\dots,P_{\ell}$ are monomials in $M(\AA)$. We also define
@@
Assume that $x_{1},x_{2},\dots,x_{n}$ are indeterminates appearing in the polynomial $f$. Suppose indeterminate $x_{j}$ appears in all monomials of $f$, then we say $f$ is homogeneous with respect to the indeterminate $x_{j}$. If $f$ is homogeneous with respect to every indeterminate $x_{j}$ where $1\leq j\leq n$, then we say $f$ is homogeneous.
Suppose $f=f(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})$ is a $\sigma^{m}$-GPI of $D$ with $\deg(f)=p$. Let $q$ be the number of indeterminates in $f$ that are not homogeneous. Suppose $q>0$. Without loss of generality, we assume $f$ is not homogeneous with respect to indeterminates $x_{1},x_{2},\cdots,x_{q}$. Set
@@
Since $f$ is a $\sigma^{m}$-GPI of $D$, one has $f_{1}$ and $f_{2}$ are also $\sigma^{m}$-GPIs of $D$. It is clear that $\deg(f_{1})$ and $\deg(f_{2})$ are less than or equal to $p$, and $f_{2}$ is homogeneous with respect to indeterminate $x_{1}$. Similarly, for $2\leq j\leq q$, set
@@
and $f_{2j}=f_{2j-2}-f_{2j-1}$. Then, the polynomial $f_{2q}$ is homogeneous with respect to indeterminates $x_{1},x_{2},\cdots,x_{n}$, and is also a $\sigma^{m}$-GPI of $D$. From the above reasoning, we obtain the following proposition.
@@
Let $D$ be a division ring and $\sigma$ be an anti-automorphism of $D$. If $D$ satisfies a $\sigma^{m}$-GPI, then $D$ satisfies a $\sigma^{m}$-GPI that is homogeneous.
@@
In other words, the polynomial $f$ is $\sigma^{m}$-linear if $f$ is a polynomial homogeneous of height 00.
Suppose $f$ is a homogeneous $\sigma^{m}$-GPI of $D$ with $\{x_{1},x_{2},\dots,x_{n}\}$ being the set of all indeterminates appearing in $f$. Let $p=\deg(f)$ and $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(f)$ $(1\leq j\leq k)$. We write $f=f(x_{j})$ to emphasize considering $f$ as a polynomial in the indeterminate $x_{j}$. If $ht(f)=t>0$, we choose $x_{j_{q}}\in X$ such that $x_{j_{q}}$ does not appear in $f$ and set
@@
Therefore, $\Delta_{j}f$ is also a homogeneous $\sigma^{m}$-GPI of $D$ satisfying $ht(\Delta_{j}f)\leq t-1$. Define
@@
Given $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ is a homogeneous polynomial in indeterminates $x_{1},x_{2},\dots,x_{n}$ and has height $t$. Let $P$ be a monomial in $f$ such that $ht(P)=t$. Then, $\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}f$ is a $\sigma^{m}$-linear polynomial in $\sum_{j=1}^{n}p_{j}$ indeterminates, where $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$ $(1\leq j\leq n)$.
Proof.  If $f$ has height $t=0$, then $p_{j}=1$ and $\Delta^{(0)}_{j}f=f$ is $\sigma^{m}$-linear for every $1\leq j\leq n$, thus the lemma is proved. Consider the case $t>0$ and $P$ is a monomial in $f$ such that $\sigma^{m}\text{-ht}(P)=t$. With $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$, we have
@@
Choose $j_{0}$ such that $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P)=p_{j_{0}}>1$ and $\deg_{x_{i}}(P)=p_{i}\leq 1$ in which $1\leq j_{0}<i\leq n$. Then
@@
in which $h_{u}$ is a polynomial such that $\sigma^{m}\text{-}\deg_{x_{n+1}}(h_{u})=u$ and $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(h_{u})=p_{j_{0}}-u$. Each polynomial $h_{u}$ is a sum of $C_{p_{j_{0}}}^{u}$ monomials $P_{u_{k}}$, where each monomial $P_{u_{k}}$ maps to $P$ under the projection $x_{n+1}\to x_{j_{0}}$ and $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P_{u_{k}})=p_{j_{0}}-u$. Hence,
@@
$\Delta_{{j_{0}}}^{p_{j_{0}}-1}P=\Delta^{(p_{j_{0}}-2)}_{{j_{0}}}h_{1}=\Delta^{(p_{j_{0}}-2)}_{{j_{0}}}\left(\sum^{C^{1}_{p_{j_{0}}}}_{k=1}P_{1_{k}}\right)=\sum^{C^{1}_{p_{j_{0}}}}_{k=1}\Delta^{(p_{j_{0}}-2)}_{j_{0}}P_{1_{k}}.$
For each $1\leq k\leq C^{1}_{p_{j_{0}}},$ we have
$\Delta^{(p_{j_{0}}-2)}_{{j_{0}}}P_{1_{k}}=\Delta^{(p_{j_{0}}-3)}_{j_{0}}\left(\Delta^{(1)}_{j_{0}}P_{1_{k}}\right)=\Delta^{(p_{j_{0}}-3)}_{j_{0}}\left(\sum^{p_{j_{0}}-2}_{u_{k}=1}h^{{}^{\prime}}_{u_{k}}\right)=\sum^{p_{j_{0}}-2}_{u_{k}=1}\Delta^{(p_{j_{0}}-3)}_{j_{0}}h^{{}^{\prime}}_{u_{k}}.$
On the other hand, each polynomial $h^{{}^{\prime}}_{u_{k}}$ has $\sigma^{m}$-$\deg_{x_{n+2}}(h^{{}^{\prime}}_{u_{k}})=u_{k}$ and $\sigma^{m}$-$\deg_{x_{j_{0}}}(h^{{}^{\prime}}_{u_{k}})=p_{j_{0}}-1-u_{k}$. Note that $h^{{}^{\prime}}_{u_{k}}$ has $C^{u_{k}}_{p_{j_{0}}-1}$ monomials, each of which maps to $P$ under the projection $x_{j_{k}}\mapsto x_{j_{0}}$ (where $j_{k}=n+1,n+2$). Similarly, for each $u_{k}>1$, we also have $\Delta_{j}^{(p_{j}-3)}h^{{}^{\prime}}_{u_{k}}=0$. Therefore, $\Delta_{j_{0}}^{(p_{j_{0}}-2)}P_{1_{k}}=\Delta_{j_{0}}^{(p_{j}-3)}h^{{}^{\prime}}_{1}$. Thus, $\Delta_{j_{0}}^{(p_{j_{0}}-1)}P=\sum_{u=1}^{C^{p_{j_{0}}}_{1}}\sum^{C^{p_{j_{0}}-1}}_{v=1}\Delta_{{j_{0}}}^{p_{j_{0}}-3}h^{"}_{uv}$, where $h"_{uv}$ are monomials. Continuing these steps, we see that $\Delta_{j}^{p_{j}-1}P$ has
@@
for $e\in\{n+1,n+2,\cdots,n+(p_{j_{0}}-1)\}$. After decomposing $\Delta_{j_{0}}^{(p_{j_{0}}-1)}P$ into monomials as above, we continue this technique for $\Delta_{j_{0}-1}^{(p_{j_{0}-1}-1)}\Delta_{j_{0}}^{(p_{j_{0}}-1)}P$, and further expand $\Delta_{j_{0}-2}^{(p_{j_{0}-2}-1)}\Delta_{j_{0}-1}^{(p_{j_{0}-1}-1)}\Delta_{j_{0}}^{(p_{j_{0}}-1)}P$. Eventually, we obtain $\Delta^{(p_{1}-1)}_{{1}}\cdots\Delta^{(p_{n}-1)}_{n}P$ as a linear polynomial in $\sum^{n}_{j=1}p_{j}$ indeterminates. Moreover, these monomials all map to $P$ under the projection satisfying the property: for every $1\leq j\leq n$, the projection of $x_{e}$ to $x_{j}$ is such that $j+\sum_{u=j+1}^{n}p_{u}\leq e\leq j+\sum_{u=j}^{n}p_{u}$.
@@
If $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}P_{i_{0}}\neq 0$ then $p_{j}^{{}^{\prime}}=\sigma^{m}$-$\deg_{x_{j}}(P_{i_{0}})\geq p_{j},$ with $1\leq j\leq n.$ However, observe that $\sum(p_{j}-1)=t\geq\sum(p^{{}^{\prime}}_{j}-1)$, which implies that $p^{{}^{\prime}}_{j}=p_{j}$ for $1\leq j\leq n.$ Using a similar method as for $P$, we have that $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}P_{i_{0}}$ is a multilinear polynomial in $\sum p_{j}$ indeterminates just like $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{n}P$. Hence, $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}f$ is a homogeneous $\sigma^{m}$-linear polynomial.
@@
If the division ring $D$ satisfies an $\sigma^{m}$-GPI of degree $n$, then there exists an $\sigma^{m}$-GPI of $D$ which is homogeneous $\sigma^{m}$-linear of degree not exceeding $n$.
@@
Proof.  By Corollary 2.3, we assume that $f$ is homogeneous $\sigma^{m}$-linear. By the hypothesis, $f$ can be represented as
@@
$=\phi\big{(}\sum_{j=1}^{n_{0}}a_{0j}\otimes\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0j}-\sum_{j=1}^{n_{0}}a_{mn_{m}}ra_{0j}\otimes b_{0j}\big{)},$
@@
00 0 $=f(r^{\sigma},\cdots,r^{\sigma^{m}})$

### v2

@@ 1. Introduction and the main result
A classic result states that if $D$ satisfies the polynomial identity, then $D$ is centrally finite (see [11]). To present this result and the extended results we aim to prove, we recall some concepts. Let $X=\{x_{1},x_{2},\cdots,x_{n}\}$ be a set of $n$ non-commutative indeterminates and $K$ be a field. We denote $K\langle X\rangle$ as the polynomial ring in indeterminates $X$ over $K$, and $R$ as a ring whose center contains $K$. A polynomial $f(X)\in K\langle X\rangle\backslash\{0\}$ is called a polynomial identity of $R$ if for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in R^{n}$, we have $f(a_{1},a_{2},\dots,a_{n})=0$. In this case, we say $R$ satisfies the polynomial identity $f$, abbreviated as PI. If $R$ satisfies some polynomial identity, then we say $R$ satisfies a polynomial identity. Since we only work with division rings, we directly assume $R=D$ is a division ring and $K$ coincides with the center $F$ of $D$.
@@
The most classical result concerning division rings satisfying polynomial identities relates to its finite-dimensionality. Specifically, it is proven in [11] that if $D$ is a division ring with center $F$ and satisfies some polynomial identity over $F$, then $D$ is centrally finite. This result has been subsequently extended in many different directions (e.g., see [1, 2, 3, 4, 5, 7, 12, 13, 14, 15]).
We now recall the basic notation used throughout. We denote $D\langle X\rangle=D*_{F}F\langle X\rangle$ as the free product of $D$ with $F\langle X\rangle$ over $F$. Each element in $D\langle X\rangle\backslash\{0\}$ is called a generalized polynomial over $D$. Furthermore, since the ring of generalized polynomials $D\langle X\rangle$ is fir, the maximal right ring of quotients of $D\langle X\rangle$ is a division ring (e.g., see [8]). This division ring is denoted as $D(X)$. Each nonzero element in $D(X)$ is called a generalized rational polynomial over $D$. Suppose $f(X)$ is a generalized rational polynomial (respectively, a generalized polynomial) over $D$. If for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in D^{n}$ such that when $f(a_{1},a_{2},\dots,a_{n})$ is defined, one has $f(a_{1},a_{2},\dots,a_{n})=0$, then we say $D$ satisfies the generalized rational identity, abbreviated as GRI (respectively, the generalized polynomial identity, abbreviated GPI) $f$, or $f$ is a generalized rational identity (respectively, a generalized polynomial identity) of $D$.
@@
In 1966, Amitsur proved that if the center $F$ is infinite and the division ring $D$ satisfies a generalized rational identity, then $D$ is centrally finite (see [2]). There are several extensions of this result. For example, in 1996, Chiba proved that if the center $F$ is infinite and the multiplicative group $D\{$$^{*}=D\backslash\{0\}$ contains a non-central proper subnormal subgroup satisfying a generalized rational identity, then $D$ is centrally finite (see [6]). Another direction is to consider identities corresponding to an anti-automorphism. Recall that a bijection $\sigma:D\to D,x\mapsto x^{\sigma},$ is called an anti-automorphism of $D$ if $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma}$ and $(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ for all $a,b\in D$. In the case where the order of $\sigma$ is $2$, $\sigma$ is called an involution of $D$. Let $m$ be a positive integer such that $\sigma^{i}\neq\mathrm{Id}$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, denote $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ as the set of indeterminates indexed corresponding to each $\sigma^{i}$ and $X_{m}=\cup_{i=0}^{m}X^{\sigma^{i}}$. We denote $D(X_{m})$ as the division ring as mentioned above. Let $f(X_{m})$ be a nonzero element in $D(X_{m})$. If for every $n$-tuple $(r_{1},r_{2},\cdots,r_{n})\in D^{n}$ such that when
@@
then we say $f(X_{m})$ is a generalized rational identity of $D$ with respect to $\sigma^{m}$. In this case, we simply say that $D$ satisfies a $\sigma^{m}$-GRI $f$ or $f$ is a $\sigma^{m}$-GRI of $D$. The ”best” result is related to when $\sigma$ is an involution. In this case, $m=1$ and $\sigma$ will be denoted as $\star$. In more detail, it is proved that if the center $F$ is infinite and $D$ satisfies a $\star$-GRI, then $D$ is centrally finite (see [9]). The case of generalized polynomial identities with respect $\sigma$ receives even more attention (e.g., see the book [3] for a survey and the paper [18] for some additional interesting results on this topic). The objective of this paper is to prove the following result.
@@ 2. Preliminaries
In short and simple terms, we prove the main result of this paper by induction on $n$, the number of indeterminates in the identity. In this section, we prove the case for $n=1$. To do it, we need to review some basic concepts and establish notation. Many results in this section may have proofs similar to those found in some reference documents, but none of the results are directly quoted from them. Therefore, for the convenience of the readers, we will provide proofs for the results if necessary. The main result we obtained in this section is Lemma  2.4, which states that if a division ring satisfies a blended linear identity in one variable, then this division ring is centrally finite.
@@
We also define the $\sigma^{m}$-height of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}\text{-}ht_{x_{j}}(P)$, as the maximum value between $(\sigma^{m}\text{-}\deg_{x_{j}}(P)-1)$ and 0. Similarly, the height of $P$, denoted $ht(P)$, is defined as
@@
[…]
@@
Assume that $x_{1},x_{2},\dots,x_{n}$ are indeterminates appearing in the polynomial $f$. Suppose indeterminate $x_{j}$ appears in all monomials of $f$, then we say $f$ is blended with respect to the indeterminate $x_{j}$. If $f$ is blended with respect to every indeterminate $x_{j}$ where $1\leq j\leq n$, then we say $f$ is blended.
Suppose $f=f(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})$ is a $\sigma^{m}$-GPI of $D$ with $\deg(f)=p$. Let $q$ be the number of indeterminates in $f$ that are not blended. Suppose $q>0$. Without loss of generality, we assume $f$ is not blended with respect to indeterminates $x_{1},x_{2},\cdots,x_{q}$. Set
@@
Since $f$ is a $\sigma^{m}$-GPI of $D$, one has $f_{1}$ and $f_{2}$ are also $\sigma^{m}$-GPIs of $D$. It is clear that $\deg(f_{1})$ and $\deg(f_{2})$ are less than or equal to $p$, and $f_{2}$ is blended with respect to indeterminate $x_{1}$. Similarly, for $2\leq j\leq q$, set
@@
and $f_{2j}=f_{2j-2}-f_{2j-1}$. Then, the polynomial $f_{2q}$ is blended and is also a $\sigma^{m}$-GPI of $D$. From the above reasoning, we obtain the following proposition.
@@
Let $D$ be a division ring and $\sigma$ be an anti-automorphism of $D$. If $D$ satisfies a $\sigma^{m}$-GPI, then $D$ satisfies a $\sigma^{m}$-GPI that is blended.
@@
In other words, the polynomial $f$ is $\sigma^{m}$-linear if $f$ is a polynomial blended of height 0.
Suppose $f$ is a blended $\sigma^{m}$-GPI of $D$ with $\{x_{1},x_{2},\dots,x_{n}\}$ being the set of all indeterminates appearing in $f$. Let $p=\deg(f)$ and $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(f)$ $(1\leq j\leq k)$. We write $f=f(x_{j})$ to emphasize considering $f$ as a polynomial in the indeterminate $x_{j}$. If $ht(f)=t>0$, we choose $x_{j_{q}}\in X$ such that $x_{j_{q}}$ does not appear in $f$ and set
@@
Therefore, $\Delta_{j}f$ is also a blended $\sigma^{m}$-GPI of $D$ satisfying $ht(\Delta_{j}f)\leq t-1$. Define
@@
Given $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ is a blended polynomial in indeterminates $x_{1},x_{2},\dots,x_{n}$ and has height $t$. Let $P$ be a monomial in $f$ such that $ht(P)=t$. Then, $\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}f$ is a $\sigma^{m}$-linear polynomial in $\sum_{j=1}^{n}p_{j}$ indeterminates, where $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$ $(1\leq j\leq n)$.
Proof.  If $f$ has height $t=0$, then $p_{j}=1$ and $\Delta_{j}f=f$ is $\sigma^{m}$-linear for every $1\leq j\leq n$, thus the lemma is proved. Consider the case $t>0$ and $P$ is a monomial in $f$ such that $\sigma^{m}\text{-ht}(P)=t$. With $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$, we have
@@
Choose $j_{0}$ such that $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P)=p_{j_{0}}>1$ and $\deg_{x_{i}}(P)=p_{i}=1$ in which $1\leq j_{0}<i\leq n$. Then
@@
in which $h_{u}$ is a polynomial such that
$\sigma^{m}\text{-}\deg_{x_{n+1}}(h_{u})=u\text{ and }\sigma^{m}\text{-}\deg_{x_{j_{0}}}(h_{u})=p_{j_{0}}-u.$
Each polynomial $h_{u}$ is a sum of $C_{p_{j_{0}}}^{u}$ monomials $P_{u_{k}}$, where each monomial $P_{u_{k}}$ maps to $P$ under the projection $x_{n+1}\to x_{j_{0}}$ and $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P_{u_{k}})=p_{j_{0}}-u$. Hence,
@@
$\Delta_{{j_{0}}}^{p_{j_{0}}-1}P=\Delta^{(p_{j_{0}}-2)}_{{j_{0}}}h_{1}.$
By induction, we obtain that $\Delta_{j}^{p_{j}-1}P$ has
@@
for $e\in\{n+1,n+2,\cdots,n+(p_{j_{0}}-1)\}$.
The remaining part of the lemma can be established by induction on $n$. Finally, we obtain $\Delta^{(p_{1}-1)}_{{1}}\cdots\Delta^{(p_{n}-1)}_{n}P$ as a linear polynomial in $\sum^{n}_{j=1}p_{j}$ indeterminates. Moreover, these monomials all map to $P$ under the projection satisfying the property: for every $1\leq j\leq n$, the projection of $x_{e}$ to $x_{j}$ is such that
$j+\sum_{u=j+1}^{n}p_{u}\leq e\leq j+\sum_{u=j}^{n}p_{u}.$
@@
If $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}P_{i_{0}}\neq 0$ then $p_{j}^{{}^{\prime}}=\sigma^{m}$-$\deg_{x_{j}}(P_{i_{0}})\geq p_{j},$ with $1\leq j\leq n.$ However, observe that $\sum(p_{j}-1)=t\geq\sum(p^{{}^{\prime}}_{j}-1)$, which implies that $p^{{}^{\prime}}_{j}=p_{j}$ for $1\leq j\leq n.$ Using a similar method as for $P$, we have that $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}P_{i_{0}}$ is a multilinear polynomial in $\sum p_{j}$ indeterminates just like $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{n}P$. Hence, $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}f$ is a blended $\sigma^{m}$-linear polynomial.
@@
If the division ring $D$ satisfies an $\sigma^{m}$-GPI of degree $n$, then there exists an $\sigma^{m}$-GPI of $D$ which is blended $\sigma^{m}$-linear of degree not exceeding $n$.
@@
Proof.  By Corollary 2.3, we assume that $f$ is blended $\sigma^{m}$-linear. By the hypothesis, $f$ can be represented as
@@
$\phi\big{(}\sum_{j=1}^{n_{0}}a_{0j}\otimes\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0j}-\sum_{j=1}^{n_{0}}a_{mn_{m}}ra_{0j}\otimes b_{0j}\big{)},$
@@
$0$ $=f(r^{\sigma},\cdots,r^{\sigma^{m}})$